OpenStudy (anonymous):

Use mathematical induction to prove that the statement is true for every positive integer n. Show your work. 2 is a factor of n2 - n + 2

OpenStudy (solomonzelman):

I will define: \(\color{#000000}{ \displaystyle f(n)=n^2-n=n(n-1) }\) Write even positive integers in a form of \(\color{#000000}{ \displaystyle 2k}\), and odd positive integers in a form of \(\color{#000000}{ \displaystyle 2k+1}\).\(\\[0.5em]\) (where \(\color{#000000}{ \displaystyle k}\) is also some/any positive integer) \(\\[0.9em]\) When you plug \(\color{#000000}{ \displaystyle 2k}\), you get: \(\color{#000000}{ \displaystyle f(2k)=2k(2k-1) }\) When you plug \(\color{#000000}{ \displaystyle 2k+1}\), you get: \(\color{#000000}{ \displaystyle f(2k+1)=2k(2k+1) }\) So, regardless of whether \(n\) is odd or even, when you plug in (any) \(n\), your output will be always divisible by 2. That's because, for any \(n\) - whether even or odd, your expression \(f(n)\) is always a product of an even×odd (where \(2k\) is an even).